Prove or disprove: If G is a finite group and some element of G has order equal

to the size of G, then G is cyclic.


1
Expert's answer
2021-07-26T09:48:04-0400

Solution:

Proof;

Let g"\epsilon" G have an order n="\#" (G).

For each i with 1"\leq" i"<" n we have gi"\ne" e,the identity of G.

The claim herein is that G="\lang g\rang"

For this ,there are n elements of {gi:0"\le" i<n}

If gi=gj for some j<i, multiply both sides of the equation by g-i=gi(-1)

We have, e=gj-i even though 1"\leq" j-i"<" n

Contrary to the above observations.

Hence G="\lang g\rang" is cyclic.